OpenStudy (anonymous):

i cant figure out x^4-52x^2+576=0

OpenStudy (anonymous):

(x^2 - 16)(x^2 - 36) (x - 4)(x + 4)(x - 6)(x + 6) x = -6, -4, 4, 6

OpenStudy (anonymous):

thanks need help with 4 more x^4-7x^2+12-0

OpenStudy (anonymous):

four more?? have you tried them?

OpenStudy (anonymous):

yes ive tried and tried all night its too hard

OpenStudy (anonymous):

it's never too hard x^4-7x^2+12 = 0 (x^2 - 3)(x^2 - 4) (x^2 - 3)(x + 2)(x - 2) \[x = \pm \sqrt3, \pm 2\]

OpenStudy (anonymous):

i dont no how to post that symbol on my pc

OpenStudy (anonymous):

just do -2, 2, -sqrt 3, sqrt 3 i'm on a pc too

OpenStudy (anonymous):

thanks again i wish i had a teacher for this

OpenStudy (anonymous):

15x^4-14x^2+3=0